OpenStudy (anonymous):

Parking in a student lot costs $2 for the first half hour and $1.75 for each hour thereafter. A partial hour is charged the same as a full hour. What is the longest time that a student can park in this lost for $7?

OpenStudy (anonymous):

Well start off by subtracting! So 7-2 = 5 dollars. So first half hour and we have 5 dollars left. Next: 5-1.75-1.75 = 1.5 So we got charged for two more hours and have nothing left to pay for either a partial or full hour. So we add it up .5h+1h+1h = 2.5H longest time to park.

OpenStudy (anonymous):

thanks, now i get it